OpenStudy (anonymous):

how do you show your work for the equation x+7=12?

hero (hero):

\[x+7=12\]Subtract 7 from both sides. \[x+7-7=12-7\]

hero (hero):

Wow, using this on a tablet is impossible.

OpenStudy (anonymous):

x + 7 =12. Subtract 7 on both sides. X - 7 -7 = 12 -7. The 7's cancel out and 12 -7 equals 5, so the answer would be: x= 5

OpenStudy (anonymous):

there should be no work to show you know right away that \(5+7=12\)